OpenStudy (pphalke):

You need to rent a moving truck for a day. You have identified two companies that rent trucks. Company A charges $45 per day plus $0.20 per mile. Company B charges $40 per day plus $0.30 per mile. For how many miles will the cost of renting be the same?

OpenStudy (yinshy):

what do you think?

OpenStudy (pphalke):

that's what I don't know.

OpenStudy (yinshy):

its fairly simple

OpenStudy (yinshy):

@elitedamosh help this person

OpenStudy (anonymous):

You can assume the miles X . Then make an equation which is 45+0.20X=40+0.30X And solve the equation by finding X value .

OpenStudy (pphalke):

how do I solve do I divide by x on each one or?

OpenStudy (pphalke):

@hinamori

OpenStudy (anonymous):

Yes divide the equation put X on a side and the real numbers on the other side with changing from positive to negative or negative to positive as you transport each one to the other side.

OpenStudy (pphalke):

so like what? I don't know how

OpenStudy (anonymous):

Ok , here's how to do it : 45+0.20X=40+0.30X 45-40=0.30X-0.20X 5=0.10X X=5/0.10 X=50 miles.

OpenStudy (anonymous):

company A: cost = .35m + 55 company B: cost = .4m + 50 solve .4m+50 = .35m+55 .05m = 5 m = 5/.05 = 100